# ALMA chemical survey of disk-outflow sources in Taurus (ALMA-DOT)

### v2

@@ 3.3 Velocity channel maps
A comprehensive atlas of the channel maps of the CO (2-1) emission is given in Appendix A. The analysis of the individual channels reveals a faint but very rich flow substructure, not visible in the moment-0 and moment-1 maps. The emission in the blue lobe is barely detectable, but the signal is strong in the red-shifted lobe, allowing for detailed inspection. A selection of channel maps illustrating the configuration of the inner flow in the red-shifted SW lobe is presented in Fig. 3. The maps show an apparent conical structure, with aperture decreasing with velocity, from about 80 at 10 km sto 55 at 26 km s(see Appendix A).
Inscribed in the conical apparent shape, one finds a series of ‘bubbles’ and ‘arcs’ of different sizes, positions and brightness. The arcs appear to be part of concatenated closed curves that can be visually fitted by ellipses, as illustrated in Fig. 4. About ten such structures can be identified, grouped in four families that we label as A, B, C, D. Each feature continuously changes size and position with channel maps and acquires progressively larger extension and separation from the star as VLSR velocity increases, but maintains its shape unchanged. Arcs ’A’ are observed at low velocity (10-16 km s, see Fig. 4a-b) and at small separation. Despite the blending of the emission on the source side and the fragmented appearance at larger distances, these structures clearly form closed bubble-like curves expanding with increasing velocities. Arcs B1 and B2 have a large extension and only show the convex branch on the side of the source. They can be followed over a wide range of velocities, from 11 to 19 km s, in which they continuously change position from 3 to 10 from the source, as illustrated in Fig. 17. Arcs C are visible only at high velocity (20-29 km s, see Fig. 4c-d. They are relatively faint, but their positional shift with velocity can be clearly recognized. Finally, the bright arc D is visible at a small distance from the source (about 09) and at intermediate velocities (20 km s, see Fig. 4c. The orientation of the features deviates slightly from that of the axis of the system. This is illustrated, for example, in panel b of Fig. 4.
@@ 3.4 Parallel Position-Velocity diagram
Figure 5 illustrates the Position-Velocity diagram obtained with a virtual slit 03 wide (comparable to the beam size) positioned along the axis (PA=228∘). The diagram (hereafter referred to as PV∥) collects emission from both circumstellar material and extended wind. The emission within offset 2 arcseconds from the source is due partly to the disk and partly from the base of the wind, whose large opening produces contributions in both red-shifted and blue-shifted velocities on each side (e.g., Klaassen et al. 2016). The emission at larger offsets and velocities comes only from the extended wind.
@@
The striking characteristic of the diagram is the fact that the emission in the wind region is not continuous, but it is distributed in a fan of separated traces on both sides of the source (superposed to a low intensity more homogeneous emission in the region and VLSR km s). In the blue lobe, at least three linear traces are seen to propagate towards larger blue-shifted VLSR velocity. On the SW red-shifted side, more than ten traces can be recognized. Within the first 4″a fan of almost linear traces opens from the source reaching high velocities, up to at least VLSR 28 km s. For each of them, the velocity appears to increase linearly with offset, in a way reminiscent of a ’Hubble law’ trend. This group of traces displays the steepest slope, that is, the highest apparent acceleration, among the features in the diagram. A second fainter fan of almost linear segments is observed between 4 and 11″, with lower VLSR velocities up to 15 km s. Interposed between the two groups, one finds two extended traces of curved shape, which reach velocities as high as 20 km s.
@@
The substructures detected in the diagram suggest a common origin with the arc- and bubble-shaped features identified in the channel maps. To illustrate this hypothesis, for each of the ellipses described in Section 3.3, and for each velocity channel, we report on the red-shifted portion of the PV∥ diagram the offset of the intersections of the fitted ellipses with the flow axis (see Fig. 4). The subscripts ’’ and ’’ indicate the intersections of the side of the ellipse closer and farther from the source, respectively. The broken lines obtained by connecting the points relative to the same side of each ellipse are strikingly coincident with sections of the emission branches in the diagram.
@@
The first group of traces reaching high velocities corresponds to arcs D (orange), C (red) and to the intersections of type ’’ of features A (green). The second group corresponds to the SW sides of the A ellipses, marked as ’’ (green). Given the inclination angle of the wind in space (=226∘.5), the A0 traces appear to correspond to the rear side of the flow and the A1 traces to the front side. Note that the A0 and A1 traces appear in the same order from the source, in a way that is consistent with the fact that in the channel maps the ellipses of the A family are concatenated rather than nested. The long curved traces correspond to the features of the B family (in blue). These filaments well reproduce the consistent shift in position with velocity illustrated in Fig. 17. The emission peak at offset 14 at a constant low velocity named ’E’ corresponds to the ridge of emission south-west to the source and almost perpendicular to the flow direction seen in Fig. 1.
@@ 3.5 Transverse Position-Velocity diagrams
Appendix B presents an atlas of the transverse Position-Velocity diagrams (hereafter PV⟂) obtained for the red lobe with a 03 wide virtual slit centered on the axis and oriented at PA=318° (i.e. perpendicular to the axis and with positive offsets toward NW). The diagrams are formed at the separation of 35 to 10″ from the source, in steps 05. Illustrative examples are given in Fig. 6, for selected separations from the star.
In all panels, the emission presents a bell shape around the axis, narrower toward high velocities. Progressing with separation from the source, the bell becomes wider, but the lateral ’walls’ maintain coherence. Outside of the bell, only low-velocity material (7.5 - 9.5 km s) is present.
@@
An evident characteristic of the structures identified in the transverse diagrams is the skew with respect to the horizontal. This property is related to the rotation of the flow. The two opposite borders of the wind, in fact, present a velocity shift due to the projection of the rotation (i.e. toroidal) velocity along the line of sight. An illustration of this effect is offered in de Valon et al. (2022), Fig. 2. The observed features display higher red-shifted velocities on the NW side (seen more clearly for the curves of the B family) consistent with a clockwise rotation sense looking from the tip of the red lobe toward the source. This sense of rotation matches that of the disk (Garufi et al., 2021).
@@
4 Interpretation of the substructures in terms of nested gas shells.
@@
In the real case, rotation and inclination introduce deformations of the traces in the projected maps. The skew of the features due to flow rotation has been described in the previous Section. On the other hand, inclination with respect to the line of sight produces a deformation of the closed curves with respect to the pure ellipse shape, as described, e.g., in De Valon (2021). We adopt for the inclination angle of the red-shifted flow with respect to the line of sight a value 46.5° . This compares well with the inclination angle of the red-shifted axial atomic jet, , determined by combining the tangential velocity measured for the most coherent [SII] emission knots on the SW side of the source in Anglada et al. (2007) with the average radial velocity measured for the same jet lobe by Mundt et al. (1990). Due to this inclination, the low-velocity part of the ellipses is expected to be flatter than the high-velocity part. However, this deformation is not easily recognizable in our maps, since the structures centered on moderate velocities are weak and/or do not have a visible upper part, while those centered on low velocities have their lower part possibly hidden by the absorption of CO gas with velocities close to Vsys.
In any case, the overall appearance of the data supports an interpretation of the observed substructures in terms of distinct rotating shells nested in physical space. As mentioned in the Introduction, the same hypothesis has been put forward for other molecular winds that exhibit similar spectro-imaging properties in high-resolution data.
However, it is important to note that HL Tau is the only object among the known cases with ring / gap substructures detected in the disk. In this case, therefore, it is possible to study the relationship between the substructures observed in the wind and those visible in the disk.
4.1 Tomographic reconstruction of the wind shells.
@@
6.5 | -5.52 | 5.03 | 11.50 | 13.90 | 1.3 | 7.8 | 8.1 | 1.65 | 17.5 | 8.5 | 12.9 | 85 | 3.2 | |
7.0 | -5.70 | 5.52 | 11.60 | 14.10 | 1.4 | 8.3 | 8.4 | 1.72 | 17.5 | 8.8 | 14.2 | 89 | 3.5 | |
7.5 | -5.89 | 5.65 | 12.08 | 14.43 | 1.5 | 8.5 | 8.9 | 1.62 | 17.5 | 9.4 | 13.8 | 83 | 3.5 | |
8.0 | -6.18 | 5.73 | 12.29 | 14.57 | 1.6 | 8.8 | 9.2 | 1.57 | 17.5 | 9.6 | 13.8 | 80 | 3.6 | |
8.5 | -6.47 | 5.84 | 12.33 | 14.67 | 1.7 | 9.1 | 9.3 | 1.61 | 17.5 | 9.8 | 14.6 | 84 | 3.7 | |
9.0 | -6.56 | 6.09 | 12.41 | 14.70 | 1.8 | 9.3 | 9.4 | 1.58 | 17.5 | 9.8 | 14.7 | 84 | 3.7 | |
9.5 | -7.10 | 6.39 | 12.40 | 14.80 | 1.9 | 9.9 | 9.4 | 1.65 | 17.5 | 9.9 | 16.4 | 92 | 4.0 | |
10 | -7.16 | 6.53 | 12.45 | 14.87 | 2.0 | 10.1 | 9.5 | 1.67 | 17.5 | 10.0 | 16.8 | 93 | 4.1 |
The series of PV⟂ diagrams illustrated in Appendix B allows, in principle, a ’tomographic reconstruction’ of the wind shells that provides information on their morphology and kinematics in the physical space. This can be done by measuring the width, inclination, and position of each detected feature in adjacent diagrams.
In practice, for each PV⟂ diagram one has to identify the SE and NW apexes of the identified features and measure the corresponding position and VLSR velocity. These coordinates are sufficient to determine the local shell radius , the toroidal velocity and the axial velocity (referring to a co-moving cylindrical coordinate system centered on the star, with directed toward PA=228°with an inclination angle =46.5° ). However, the task is obstructed by the confusion generated by the faintness of the signal and the overlap of structures at low VLSR. The main difficulty lies in identifying a coherent feature that can be followed in its evolution from a diagram to the next. However, we illustrate the measurement by focusing on the features of family B, which are more clearly identified and have a regular shape. Note that a close inspection of the datacube projections reveals that the B1 feature is actually composed of two intersecting curves that we label B1a and B1b. The analysis of the other fainter arcs, requiring a dedicated procedure, will be presented in the next paper of the series (Nony et al., in prep).
To ensure the identification of the apexes in a way that is as reliable as possible, a cross-check was first performed with the positions of the emission peaks in common with the other projections of the datacube. Secondly, a mathematical fit of the sequence of peaks along the identified traces was performed, assuming an ellipse as the geometrical curve closest to the shape of the features. The SE and NW extremes of the fitted ellipses, that is, the points on the curve furthest from the axis, are considered the vertices of each feature at the shell’s edges, and their coordinates are recorded. The procedure is illustrated in Fig. 8, and details on sampling criteria and fitting routine, as well as a discussion of inherent limitations, are available in Appendix B.
@@
where +7.1 km sand =46.5∘.444 In writing these formulas we neglect flow wiggling and precession. A check on the validity of this assumption is reported in Appendix C. In addition, the tangent angle to each shell is found with a linear fit to the variation of with . If the wind flows along the shell, the poloidal velocity is tangential to the surface, and . To validate this assumption, one should measure independently . Unfortunately, this measurement is complicated by the inclination of the flow with respect to the l.o.s., which distributes the information at displaced offsets (see Fig. 21), in regions where the emission traces are faint and/or confused. This makes the test conditions prohibitive in almost all cases. However, we could measure and confirm the validity of the hypothesis at least for a few positions, as described in Appendix D. In the following, it is assumed that the flow streams along the shell surfaces.
A similar but not identical tomographic analysis has been performed for the study of the flow from the young star DG Tau B in de Valon et al. (2022). The comparison of the two approaches is discussed in Appendix E.
@@
The parameters derived from the fit for the offsets in which the B substructures are well identified are given in Footnote 3. The same quantities are illustrated in Fig. 10 as a function of both the offset and the deprojected distance from the source along the flow axis. For a visualization of the position of the three considered shells in the flow portion examined, see Fig. 9, which reports the spatial offsets of the SE and NW apexes of the ellipse fit ( and in Footnote 3) superposed to a moment 0 intensity map integrated over the velocity range 10.6 - 16.2 km s, (characteristic of the studied traces), rotated to have the axis along the vertical.
@@
The three shells turn out to have a similar trend for the radius, that increases from about 5102 to 103 au for distances from 103 to 2103 au from the star, and appear indeed to be nested, with B1a the innermost one and B2 the external layer. The smaller opening angle of B2 with respect to the one determined for B1a and B1b at offsets closer to the source, may indicate recollimation at large distances. The axial velocity , between 7 and 11 km sat the distances sampled, is higher for the inner shells and increases monotonically. In shell B2 the increase is less pronounced after 1.6103 au, and the same occurs for B1a and B1b, but closer to the source, after 1.3103 au. The poloidal velocity , proportional to , shows a similar trend. The toroidal velocity is constant on average along a single shell, but is progressively higher toward the axis, with an average of 1.85, 1.77 and 1.64 km s, for B1a, B1b, and B2, respectively. The specific angular momentum varies between 950 and 1700 au km s. Its value increases monotonically and appears to be similar in the three shells. It is also interesting to estimate the dynamical time at each position along the propagation streamline based on the derived results for . An illustration of this point is reported in Appendix F.

### v3

@@ 3.3 Velocity channel maps
A comprehensive atlas of the channel maps of CO (2-1) emission is given in Appendix A. The analysis of the individual channels reveals a faint but very rich flow substructure, not visible in the moment-0 and moment-1 maps. The emission in the blue lobe is barely detectable, but the signal is strong in the redshifted lobe, thus allowing a detailed inspection. A selection of channel maps illustrating the configuration of the inner flow in the redshifted SW lobe is presented in Fig. 3. The maps show an apparent conical structure, with aperture decreasing with velocity, from about 80 at 10 km sto 55 at 26 km s(see Appendix A).
Inscribed in the conical apparent shape, one finds a series of ‘bubbles’ and ‘arcs’ of different sizes, positions and brightness. The arcs appear to be part of concatenated closed curves that can be visually fitted by ellipses, as illustrated in Fig. 4. About ten such structures can be identified, grouped in four families that we label A, B, C, and D. Each feature continuously changes size and position with channel maps and acquires progressively larger extension and separation from the star as VLSR velocity increases, but maintains its shape unchanged.
The arcs of family A are observed at low velocity (10-16 km s, see Fig. 4a-b) and at small separation. Despite the blending of the emission on the source side and the fragmented appearance at larger distances, these structures clearly form closed bubble-like curves expanding with increasing velocities. Arcs B1 and B2 have a large extension and only show the convex branch on the side of the source. They can be followed over a wide range of velocities, from 11 to 19 km s, in which they continuously change position from 3 to 10 from the source, as illustrated in Fig. 17. Arcs C are visible only at high velocity (20-29 km s, see Fig. 4c-d). They are relatively faint, but their positional shift with velocity can be clearly recognized. Finally, the bright arc D is visible at a small distance from the source (about 09) and at intermediate velocities (20 km s, see Fig. 4c).
In the following, we define as the axis of the system the symmetry axis of the disk. Its projection on the plane of the sky is shown as a solid green line in Fig. 4, and is taken to be the reference axis of the positional offset from the source, with positive offsets toward SW (PA=228 north to east). The orientation of the features deviates slightly from this direction, as is evident, for example, for the features of the B family in panel b of Fig. 4.
@@ 3.4 Parallel position-velocity diagram
Figure 5 illustrates the position-velocity diagram obtained with a virtual slit 03 wide (comparable to the beam size) positioned along the axis (PA=228∘). The diagram (hereafter referred to as PV∥) collects emission from both circumstellar material and extended wind. The emission within offset 2 arcseconds from the source is due partly to the disk and partly from the base of the wind, whose large opening produces contributions in both redshifted and blueshifted velocities on each side (e.g., Klaassen16). The emission at larger offsets and velocities comes only from the extended wind.
@@
The striking characteristic of the diagram is the fact that the emission in the wind region is not continuous, but it is distributed in a fan of separated traces on both sides of the source (superposed to a low intensity more homogeneous emission in the region and VLSR km s). In the blue lobe, at least three linear traces are seen to propagate toward larger blueshifted VLSR velocity. On the SW redshifted side, more than ten traces can be recognized. Within the first 4″a fan of almost linear traces opens from the source reaching high velocities, up to at least VLSR 28 km s. For each of them, the velocity appears to increase linearly with offset, in a way reminiscent of a Hubble law trend. This group of traces displays the steepest slope, that is, the highest apparent acceleration, among the features in the diagram. A second fainter fan of almost linear segments is observed between 4 and 11″, with lower VLSR velocities up to 15 km s. Interposed between the two groups, two extended traces of curved shape reach velocities as high as 20 km s.
@@
The substructures detected in the diagram suggest a common origin with the arc- and bubble-shaped features identified in the channel maps. To illustrate this hypothesis, for each of the ellipses described in Section 3.3, and for each velocity channel, we report on the redshifted portion of the PV∥ diagram the offset of the intersections of the fitted ellipses with the flow axis (see Fig. 4). The subscripts and indicate the intersections of the side of the ellipse closer and farther from the source, respectively. The broken lines obtained by connecting the points relative to the same side of each ellipse are strikingly coincident with sections of the emission branches in the diagram.
@@
The first group of traces reaching high velocities corresponds to arcs D (orange), C (red) and to the intersections of type of features A (green). The second group corresponds to the SW sides of the A ellipses, marked as “” (green). Given the inclination angle of the wind in space (=226∘.5), the A0 traces appear to correspond to the rear side of the flow and the A1 traces to the front side. We note that the A0 and A1 traces appear in the same order from the source, in a way that is consistent with the fact that in the channel maps the ellipses of the A family are concatenated rather than nested. The long curved traces correspond to the features of the B family (in blue). These filaments well reproduce the consistent shift in position with velocity illustrated in Fig. 17. The emission peak at offset 14 at a constant low velocity denoted E corresponds to the ridge of emission southwest of the source and almost perpendicular to the flow direction seen in Fig. 1.
@@ 3.5 Transverse position-velocity diagrams
Appendix B presents an atlas of the transverse position-velocity diagrams (hereafter PV⟂) obtained for the red lobe with a 03 wide virtual slit centered on the axis and oriented at PA=318° (i.e., perpendicular to the axis and with positive offsets toward NW). The diagrams are formed at the separation of 35 to 10″ from the source, in steps 05 (see Fig. 18). Illustrative examples are given in Fig. 6, for selected separations from the star.
In all panels, the emission presents a bell shape around the axis, narrower toward high velocities. Progressing with separation from the source, the bell becomes wider, but the lateral walls maintain coherence. Outside of the bell, only low-velocity material (7.5 - 9.5 km s) is present.
@@
An evident characteristic of the structures identified in the transverse diagrams is the skew with respect to the horizontal. This property is related to the rotation of the flow. The two opposite borders of the wind, in fact, present a velocity shift due to the projection of the rotation (i.e., toroidal) velocity along the line of sight. An illustration of this effect is offered in DeValon22, Fig. 2. The observed features display higher redshifted velocities on the NW side (seen more clearly for the curves of the B family) consistent with a clockwise rotation observed from the tip of the red lobe toward the source. This direction of rotation matches that of the disk (Garufi21).
@@
4 Interpretation of the substructures in terms of nested gas shells
@@
In the real case, rotation and inclination introduce deformations of the traces in the projected maps. The skew of the features due to flow rotation has been described in the previous Section. On the other hand, inclination with respect to the line of sight produces a deformation of the closed curves with respect to the pure ellipse shape, as described in DeValon21, among others. We adopt for the inclination angle of the redshifted flow with respect to the line of sight a value 46.5° . This compares well with the inclination angle of the redshifted axial atomic jet, , determined by combining the tangential velocity measured for the most coherent [SII] emission knots on the SW side of the source in Anglada07 with the average radial velocity measured for the same jet lobe by Mundt90. Due to this inclination, the low-velocity part of the ellipses is expected to be flatter than the high-velocity part. However, this deformation is not easily recognizable in our maps, since the structures centered on moderate velocities are weak and/or do not have a visible upper part, while those centered on low velocities have their lower part possibly hidden by the absorption of CO gas with velocities close to Vsys.
In any case, the overall appearance of the data supports an interpretation of the observed substructures in terms of distinct nested rotating shells in physical space. As mentioned in the Introduction, the same hypothesis has been put forward for other molecular winds that exhibit similar spectro-imaging properties in high-resolution data.
However, it is important to note that HL Tau is the only object among the known cases with a ring–gap configuration detected in the disk. In this case, therefore, it is possible to study the relationship between the substructures observed in the wind and those visible in the disk.
4.1 Tomographic reconstruction of the wind shells
@@
| 6.5 | -5.52 | 5.03 | 11.50 | 13.90 | 1.3 | 7.8 | 8.1 | 1.65 | 17.5 | 8.5 | 12.9 | 85 | 3.2 | |
| 7.0 | -5.70 | 5.52 | 11.60 | 14.10 | 1.4 | 8.3 | 8.4 | 1.72 | 17.5 | 8.8 | 14.2 | 89 | 3.5 | |
| 7.5 | -5.89 | 5.65 | 12.08 | 14.43 | 1.5 | 8.5 | 8.9 | 1.62 | 17.5 | 9.4 | 13.8 | 83 | 3.5 | |
| 8.0 | -6.18 | 5.73 | 12.29 | 14.57 | 1.6 | 8.8 | 9.2 | 1.57 | 17.5 | 9.6 | 13.8 | 80 | 3.6 | |
| 8.5 | -6.47 | 5.84 | 12.33 | 14.67 | 1.7 | 9.1 | 9.3 | 1.61 | 17.5 | 9.8 | 14.6 | 84 | 3.7 | |
| 9.0 | -6.56 | 6.09 | 12.41 | 14.70 | 1.8 | 9.3 | 9.4 | 1.58 | 17.5 | 9.8 | 14.7 | 84 | 3.7 | |
| 9.5 | -7.10 | 6.39 | 12.40 | 14.80 | 1.9 | 9.9 | 9.4 | 1.65 | 17.5 | 9.9 | 16.4 | 92 | 4.0 | |
| 10 | -7.16 | 6.53 | 12.45 | 14.87 | 2.0 | 10.1 | 9.5 | 1.67 | 17.5 | 10.0 | 16.8 | 93 | 4.1 |
The series of PV⟂ diagrams illustrated in Appendix B allows, in principle, a tomographic reconstruction of the wind shells that provides information on their morphology and kinematics in the physical space. This can be done by measuring the width, inclination, and position of each detected feature in adjacent diagrams.
In practice, for each PV⟂ diagram one has to identify the SE and NW apexes of the identified features and measure the corresponding position and VLSR velocity. These coordinates are sufficient to determine the local shell radius , the toroidal velocity and the axial velocity (referring to a co-moving cylindrical coordinate system centered on the star, with directed toward PA=228°with an inclination angle =46.5° ). However, the task is obstructed by the confusion generated by the faintness of the signal and the overlap of structures at low VLSR. The main difficulty lies in identifying a coherent feature that can be followed in its evolution from a diagram to the next. However, we illustrate the measurement by focusing on the features of family B, which are more clearly identified and have a regular shape. We note that a close inspection of the datacube projections revealed that the B1 feature is actually composed of two intersecting curves that we label B1a and B1b. The analysis of the other fainter arcs, requiring a dedicated procedure, will be presented in the next paper of the series (Nony et al., in prep).
To ensure identification of the apexes in a way that is as reliable as possible, a cross-check was first performed with the positions of the emission peaks in common with the other projections of the datacube. Secondly, a mathematical fit of the sequence of peaks along the identified traces was performed, assuming an ellipse as the geometrical curve closest to the shape of the features. The SE and NW extremes of the fitted ellipses, that is, the points on the curve furthest from the axis, were considered the vertices of each feature at the edges of the shell, and their coordinates were recorded. The procedure is illustrated in Fig. 8, and details on sampling criteria and fitting routine, as well as a discussion of inherent limitations, are available in Appendix B.
@@
where +7.1 km sand =46.5∘.444In writing these formulas we neglected flow wiggling and precession. The validity of this assumption is discussed in Appendix C. Additionally, the tangent angle to each shell was found with a linear fit to the variation of with . If the wind flows along the shell, the poloidal velocity is tangential to the surface, and . To validate this assumption, one should measure independently . Unfortunately, this measurement is complicated by the inclination of the flow with respect to the l.o.s., which distributes the information at displaced offsets (see Fig. 22), in regions where the emission traces are faint and/or confused. This makes the test conditions prohibitive in almost all cases. However, we could measure and confirm the validity of the hypothesis at least for a few positions, as described in Appendix D. In the following, it is assumed that the flow streams along the shell surfaces. A similar but not identical tomographic analysis has been performed for the study of the flow from the young star DG Tau B in DeValon22. The comparison of the two approaches is discussed in Appendix E.
4.2 Derived parameters of the wind shells
@@
The results of the tomographic analysis for the offsets in which the B substructures are well identified are given in Footnote 3. The same quantities are illustrated in Fig. 10 as a function of both the offset and the deprojected distance from the source along the flow axis. For a visualization of the position of the three considered shells in the flow portion examined, see Fig. 9, which reports the spatial offsets of the SE and NW apexes of the ellipse fit ( and in Footnote 3) superposed to a moment 0 intensity map integrated over the velocity range 10.6 - 16.2 km s(characteristic of the studied traces), rotated to have the axis along the vertical.
@@
The three shells turn out to have a similar trend for the radius, that increases from about 5102 to 103 au for distances from 103 to 2103 au from the star, and appear indeed to be nested, with B1a the innermost one and B2 the external layer. The opening angle of B2, smaller than the ones of B1a and B1b, that were determined at offsets closer to the source, may indicate recollimation at large distances. The axial velocity , between 7 and 11 km sat the distances sampled, is higher for the inner shells and increases monotonically. In shell B2 the increase is less pronounced after 1.6103 au, and the same occurs for B1a and B1b, but closer to the source, after 1.3103 au. The poloidal velocity , proportional to , shows a similar trend. The toroidal velocity is constant on average along a single shell, but is progressively higher for the shells closer to the axis, with an average along the flow of 1.85, 1.77 and 1.64 km s, for B1a, B1b, and B2, respectively. The specific angular momentum varies between 950 and 1700 au km s. Its value increases monotonically and appears to be similar in the three shells. It is also interesting to estimate the dynamical time at each position along the propagation streamline based on the derived results for . An illustration of this point is reported in Appendix F.